Database last updated on Sun, 01 Feb 2026 14:40:50 +0000 by localhost

Critical warnings from b-files' comments for days between 2024073 and 2024073 (0 sec, 1 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
hp:Spectrometer alignment trouble.11B07324.224TORON#22417:22:301-1

Instrument constant changes for the week before 2024073 (3 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 3 rows)

Non-critical warnings from b-files' comments for days between 2024073 and 2024073 (0 sec, 2 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
hg:HOME key pressed199B07324.224TORON#22417:23:121-1
pd:Going to menu199B07324.224TORON#22417:23:151-1

Measurements and tests summary.


_jday_loc_serial_skc
073TORONTO#224menu
(0 sec)

Number of days since last dome/window cleaning (0 sec, 2 rows)

serialdays
224more than a year

Missing files (0 sec, 1 rows)

Seriallast datadays ago
22420240740-1

Run/Stop results for last week. (0 sec, 1 rows)

SerialTotalFailed 0Failed darkFailed 1Failed 2Failed 3Failed 4Failed 5Failed double
2243001000000

Ozone results for good DS obs (std<2.5DU, mu<4) within 3 minutes between 2 Brewers. (0 sec)

No ozone measurements to display.

Unsorted comments (2 sec, 11932 rows)

SerialtypeTextCountYYYYJJJ
224env:DosBox12024073
224cubdsp:112024073
224fre_after:mm 8626 864012024073
224fre_after:mm 7772 863212024073
224fre_after:fr 5878 588912024073
224fre_after:hp 8153 816412024073
224fre_after:hg 6225 623612024073
224fre_after:da_lo 11811 1181712024073

The END (0 sec, 2 rows)

If you have questions and/or comments about this page please contact Volodya Savastiouk